# On a class of nonlinear BGK-type kinetic equations with density dependent collision rates

### v2

@@ 1 Introduction
The class of equations that we study are as follows,
@@
where $f:=f(t,x,v)\geq 0$, $(t,x,v)\in(0,+\infty]\times\mathbb{T}^{d}\times\mathbb{R}^{d}$, $d\geq 1$, $\rho_{f}(t,x):=\int_{\mathbb{T}^{d}}f(t,x,v)\,\mathrm{d}v$ is the spatial density, $\alpha\in\mathbb{R}$ and $\mathcal{M}(v)=(2\pi)^{-d/2}\exp(-|v|^{2}/2)$ is a Gaussian function (or the normalised Maxwellian distribution with 00 mean velocity, unit temperature and the Boltzmann constant). Here, $\mathbb{T}^{d}$ denotes the $d-$dimensional unit torus.
@@
    $\begin{split}\partial_{t}\rho(t,x)&=\nabla_{x}\cdot\left(\rho^{-\alpha}(t,x)\nabla_{x}\rho(t,x)\right),\\ \rho(0,x)&=\rho_{\text{in}}(x).\end{split}$ (1.3)
@@ 1.1 The models
The class of equations (1.1) also fall into a large class of kinetic transport equations, called the run and tumble equations in mathematical biology. The run and tumble equations, introduced in [1, 33], are used in modelling the movement of bacteria that respond to chemical gradients. The collision rate $\lambda$ in (1.4) corresponds to the tumbling rate of bacteria. Although, it is common to assume $\lambda$ to be a constant to ease the mathematical analysis, in physically more relevant cases $\lambda$ depends on the gradient of the chemoattractant density function which solves a Poisson-type equation [28]. On the other hand, the run and tumble equation differs from similar kinetic equations (1.1) in various ways, such as the confinement mechanism and non-explicit equilibrium solutions [37]. Under some conditions on the initial mass, it may exhibit finite time extinction of solutions [10], similar to its diffusive limit, Keller-Segel equation. See also [21, 22] for recent hypocoercivity results on various run and tumble equations.
@@
This equation share some similarities with the Landau equation which is a model for plasma dynamics. Equations of this type have been studied in [25, 2] as part of the program to utilise Di Giorgi methods in kinetic theory.
@@
In the subsequent sections, we justify the diffusive limit rigorously. Here, in order to better understand how this comes about we show a formal result by a Hilbert series expansion.
For the Hilbert series expansion, we formally write $f_{\varepsilon}=f_{0}+\varepsilon f_{1}+\mathcal{O}(\varepsilon^{2})$, $\rho_{f_{\varepsilon}}=\rho_{0}+\varepsilon\rho_{1}+\mathcal{O}(\varepsilon^{2})$ and substitute these expansions into (1.2). Letting $\varepsilon\rightarrow 0$ we immediately find $f_{0}=\mathcal{M}\rho_{0}$. Assuming the boundedness from below of $\rho_{0}$, we now expand the power law as a binomial series to obtain
@@ 1.3 State of the art
In a series of papers [5, 7, 6], Bardos, Golse, Perthame and Sentis study the Rosseland approximation for the radiative transfer equation which is a nonlinear kinetic transport equation similar to (1.1). This equation models the transport of photons in a starlike medium. The term corresponding to the collision rate $\lambda(\rho):=\rho^{\alpha}$ in (1.1) is the opacity of the medium $\sigma(T)$ at the internal energy $T$. In [5, 7, 6], the authors derive the so-called Rosseland equation as a hydrodynamic limit of the radiative transfer equation. The Rosseland equation is a degenerate parabolic equation similar to (1.3) with $\alpha=1$ with a constant boundary value. Particularly [7] covers partially the case corresponding to $-1<\alpha<0$ in (1.1) without requiring the monotonicity of $\sigma$. They employ Schauder’s fixed point argument, some compactness results and energy estimates to prove the well-posedness of the kinetic equation.
In [18], Dolbeault, Markovich, Oelz and Schmeiser study a different type of kinetic transport equation with a nonlinear relaxation term towards a generalised local Gibbs state. The equation is posed in $(x,v)\in(\mathbb{R}^{3}\times\mathbb{R}^{3})$ with a given confining potential. In the diffusive limit, they obtain drift-diffusion equations with a nonlinear diffusion term of porous medium type, corresponding to the case $-\frac{2}{3}<\alpha<1$ in (1.3). The drift term arises from the confining potential. They prove the existence and uniquness of the solutions of the kinetic equation for initial data bounded by equilibrium distributions. The diffusion limit is obtained by using a div-curl lemma based compactness argument; thus, is not quantitative.
More recently, Anceschi and Zhu, in [2], study the Cauchy problem and diffusion asymptotics of a nonlinear kinetic Fokker-Planck type equation (1.5) where $\alpha\in[0,1]$, and $(x,v)\in\mathbb{T}^{d}\times\mathbb{R}^{d}$. They prove well-posedness and propagation of Maxwellian lower bounds of the solutions by using some results based on Di Giorgi-Nash-Moser theory and the Harnack inequality. They derive quantitative diffusion asymptotics by combining entropic hypocoercivity, relative $\phi$-entropy, and barrier function methods. In this paper, we will follow their broad strategy for showing the quantitative diffusive limit. The key difference to our work is that the kinetic Fokker-Planck equation has smooth solutions.
@@
Hydrodynamic limits make up a large area of active research in kinetic theory, thanks to its connection to the Hilbert’s sixth problem entitled "Mathematical Treatment of the Axioms of Physics" which concerns developing rigorously the limiting processes connecting the atomistic description of the matter with the laws of motion of continua. Therefore, the most classical problem in this area is justifying the limit from the Boltzmann equation to either Euler or Navier-Stokes equations. Breakthrough results are due to Bardos, Golse and Levermore in [3, 4] and then due to Golse and Saint-Raymond in [23]. For a good review in this direction we refer the reader to [32]. In [31], Saint-Raymond provided a complete derivation of the Navier–Stokes–Fourier equations from a BGK equation (1.4) when $\lambda$ is a constant.
@@ 2.1 Main results
We present our main results with three theorems. In what follows, we denote the weighted $L^{2}$ norm of a function $f$ as $\|f\|_{L^{2}_{x,v}(\mathcal{M}^{-1})}:=\left(\int_{\mathbb{T}^{d}}\int_{\mathbb{R}^{d}}|f(t,x,v)|^{2}\mathcal{M}^{-1}(v)\,\mathrm{d}x\,\mathrm{d}v\right)^{\frac{1}{2}}$.
@@
Let $\varepsilon>0,A>1,\alpha\in\mathbb{R}$, and let $f_{\varepsilon}(t,x,v)$ be the solution to (1.2) with $\|f_{\varepsilon}\|_{L^{1}_{x,v}}=1$. Suppose that $\rho_{f_{\varepsilon}}(t,x)=\int_{\mathbb{R}^{d}}f_{\varepsilon}(t,x,v)\,\mathrm{d}v$ satisfies $A^{-1}\leq\rho_{f_{\varepsilon}}(t,x)\leq A$ for all $(t,x)\in[0,\infty)\times\mathbb{T}^{d}$. Then, there exists $\varepsilon_{0}>0$ such that if $\varepsilon\in(0,\varepsilon_{0})$ then
@@
$\gamma=\frac{1}{4CA^{\alpha}(1+A^{2\alpha})},$
and $C>0$ is a constant independent from $\varepsilon$.
@@ 2.2 Preliminary results on hydrodynamic quantities
$\begin{split}\rho_{\varepsilon}(t,x)&:=\int f_{\varepsilon}(t,x,v)\,\mathrm{d}v,\\ j_{\varepsilon}(t,x)&:=\int f_{\varepsilon}(t,x,v)v\,\mathrm{d}v,\\ E_{\varepsilon}(t,x)&:=\int f_{\varepsilon}(t,x)(v\otimes v-I)\,\mathrm{d}v.\end{split}$ (2.2)
@@
Let $f_{\varepsilon}$ be a solution to (1.2) and $\rho_{\varepsilon},j_{\varepsilon},E_{\varepsilon}$ be the hydrodynamic quantities associated to $f_{\varepsilon}$ as defined in (2.2). Then,
@@
$\partial_{t}(\varepsilon\nabla_{x}\cdot(\rho_{\varepsilon}^{-\alpha}j_{\varepsilon}))$ $=-\alpha\nabla_{x}\cdot(\rho_{\varepsilon}^{-\alpha-1}j(\nabla_{x}\cdot j_{\varepsilon}))-\nabla_{x}\cdot(\rho_{\varepsilon}^{-\alpha}\nabla_{x}\cdot E_{\varepsilon})-\nabla_{x}\cdot(\rho_{\varepsilon}^{-\alpha}\nabla_{x}\rho_{\varepsilon})-\frac{1}{\varepsilon}\nabla_{x}\cdot j_{\varepsilon}.$
@@
Similarly from equations (2.3), (2.4), we can see that we expect $-\frac{1}{\varepsilon}\rho_{\varepsilon}^{\alpha}j_{\varepsilon}-\nabla_{x}\rho_{\varepsilon}\rightarrow 0,$ as $t\rightarrow\infty$.
@@ 3 Well-posedness
for $t\in\mathbb{R}_{+},x\in\mathbb{T}^{d},v\in\mathbb{R}^{d}$ and $\lambda:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$. We assume that $\lambda$ satisfies the following bounds,
@@
$\partial_{t}r+v\cdot\nabla_{x}r+\lambda(\rho_{f})r$ $=\rho_{f}\lambda(\rho_{f})\mathcal{M},$
@@
By the method of characteristics we may explicitly solve for $r$ and use $r(t=0)\geq 0$ to show that $r\geq 0$, and so $g\leq A\mathcal{M}$. Similarly, taking $r=g-A^{-1}\mathcal{M}$, we find that $A^{-1}\mathcal{M}\leq g$ and so $g\in\mathcal{V}$.
Let us now show that $\Gamma$ is a contraction on $\mathcal{V}$. To this end, we let $f_{1},f_{2}\in\mathcal{V}$, set $g_{1}=\Gamma(f_{1}),g_{2}=\Gamma(f_{2})$ and $w=g_{1}-g_{2}$. Then $w$ solves
@@ 4 Long-time behaviour
In this section, we show that the solution to Equation (1.2) converges to equilibrium in $L^{2}(\mathcal{M}^{-1})$ as $t\rightarrow\infty$. We achieve this by adapting the hypocoercivity techniques developed in [19]. Particularly, we have to pay attention to ensure that the nonlinear terms do not cause us problems.
@@ 4.1 $L^{2}$-hypocoercivity à la Dolbeault-Mouhot-Schmeiser
We start with defining the operator $(I-\Delta_{x})^{-1}$ for $x\in\mathbb{T}^{d}$ by $\mathscr{F}((I-\Delta_{x})^{-1}u):=\frac{1}{1+2\pi|k|^{2}}\hat{u}$.
@@
$\|(I-\Delta_{x})^{-1}\partial_{x_{i}}u\|_{L^{2}_{x}}^{2}$ $\leq\|u-\bar{u}\|_{L^{2}_{x}}^{2}\leq\|u\|_{2}^{2},$ (4.2)
$\|(I-\Delta_{x})^{-1/2}\nabla_{x}u\|_{L^{2}_{x}}^{2}$ $\geq\|u-\bar{u}\|_{L^{2}_{x}}^{2}.$ (4.3)
@@
Notice that this operator commutes with taking gradients in $x$. It is also positive and self adjoint and so has a square root which is given by $\mathscr{F}((I-\Delta_{x})^{-1/2}u)=\frac{1}{(1+2\pi|k|^{2})^{1/2}}\hat{u}$. The rest follows by direct computation. Then, we have, denoting $\bar{u}=\int u(x)\,\mathrm{d}x$,
$\|(I-\Delta_{x})^{-1}\partial^{2}_{x_{i},x_{j}}u\|_{L^{2}_{x}}^{2}$ $=\sum_{k\in\mathbb{Z}^{d}}\left|\frac{2\pi k_{i}k_{j}}{1+2\pi|k|^{2}}\hat{u}\right|^{2}\leq\sum_{k\in\mathbb{Z}^{d}}|\hat{u}(k)|^{2}=\|u\|_{L^{2}_{x}}^{2},$
$\|(I-\Delta_{x})^{-1}\partial_{x_{i}}u\|_{L^{2}_{x}}^{2}$ $=\sum_{k\in\mathbb{Z}^{d}}\left|\frac{2\pi k_{i}}{1+2\pi|k|^{2}}\hat{u}\right|^{2}\leq\sum_{k\in\mathbb{Z}^{d},k\neq 0}|\hat{u}(k)|^{2}=\|u-\bar{u}\|_{L^{2}_{x}}^{2}\leq\|u\|_{L^{2}_{x}}^{2},$
$\|(I-\Delta_{x})^{-1/2}\nabla_{x}u\|_{L^{2}_{x}}^{2}$ $=\sum_{k\in\mathbb{Z}}\frac{2\pi|k|^{2}}{1+2\pi|k|^{2}}|\hat{u}(k)|^{2}\geq\sum_{k\in\mathbb{Z},k\neq 0}|\hat{u}(k)|^{2}=\|u-\bar{u}\|_{L^{2}_{x}}^{2}.$
@@
First we look at the dissipation of the $L^{2}$ norm. Occasionally we use the shorter notations $f_{\varepsilon}(v)=f_{\varepsilon}(t,x,v)$ or $f_{\varepsilon}(t)=f_{\varepsilon}(t,x,v)$. We have,
@@
$=-\frac{1}{\varepsilon^{2}}\int_{\mathbb{T}^{d}}\rho_{\varepsilon}^{\alpha}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\left(\frac{f_{\varepsilon}(u)}{\mathcal{M}(u)}-\frac{f_{\varepsilon}(v)}{\mathcal{M}(v)}\right)^{2}+\frac{f_{\varepsilon}^{2}(u)}{\mathcal{M}^{2}(u)}-\frac{f_{\varepsilon}(v)}{\mathcal{M}^{2}(v)}\right)\mathcal{M}(u)\mathcal{M}(v)\,\mathrm{d}u\,\mathrm{d}v\,\mathrm{d}x$

### v3

@@ 1 Introduction
The class of equations that we study is as follows,
@@
where $f:=f(t,x,v)\geq 0$, $(t,x,v)\in(0,+\infty]\times\mathbb{T}^{d}\times\mathbb{R}^{d}$, $d\geq 1$, $\rho_{f}(t,x):=\int_{\mathbb{R}^{d}}f(t,x,v)\,\mathrm{d}v$ is the spatial density, $\alpha\in\mathbb{R}$ and $\mathcal{M}(v)=(2\pi)^{-d/2}\exp(-|v|^{2}/2)$ is a Gaussian function (or the normalised Maxwellian distribution with 0 mean velocity, unit temperature and the Boltzmann constant). Here, $\mathbb{T}^{d}$ denotes the $d$-dimensional unit torus.
@@
    $\begin{cases}\begin{split}\partial_{t}\rho(t,x)&=\nabla_{x}\cdot\left(\rho^{-\alpha}(t,x)\nabla_{x}\rho(t,x)\right),\\ \rho(0,x)&=\rho_{\text{in}}(x).\end{split}\end{cases}$ (1.3)
@@ 1.1 The models
The class of equations (1.1) also falls into a large class of kinetic transport equations, called the run and tumble equations in mathematical biology. The run and tumble equations, introduced in [1, 35], are used in modelling the movement of bacteria that respond to chemical gradients. The collision rate $\lambda$ in (1.4) corresponds to the tumbling rate of bacteria. Although, it is common to assume $\lambda$ to be a constant to ease the mathematical analysis, in physically more relevant cases $\lambda$ depends on the gradient of the chemoattractant density function which solves a Poisson-type equation [30]. On the other hand, the run and tumble equation differs from similar kinetic equations (1.1) in various ways, such as the confinement mechanism and non-explicit equilibrium solutions [39]. Under some conditions on the initial mass, it may exhibit finite time extinction of solutions [10], similar to its diffusive limit, Keller-Segel equation. See also [23, 24] for recent hypocoercivity results on various run and tumble equations.
@@
This equation shares some similarities with the Landau equation which is a model for plasma dynamics. Equations of this type have been studied in [27, 2] as part of the program to utilise De Giorgi methods in kinetic theory.
@@
In the subsequent sections, we justify the diffusive limit rigorously. Here, in order to better understand how this comes about we perform a formal analysis of the diffusive limit by a Hilbert series expansion. That is, forgetting issues of convergence of the series, we assume that we can expand the solution $f_{\varepsilon}$ as a formal power series in $\varepsilon$ over the ring of test functions. In this purely formal expansion, each coefficient corresponds to the effect of an order of $\varepsilon$ on the solution. It turns out that only effects up to order one in $\varepsilon$ have an effect on our solution in this case.
We thus write $f_{\varepsilon}=f_{0}+\varepsilon f_{1}+\mathcal{O}(\varepsilon^{2})$ where $f_{0},f_{1}\in C^{\infty}_{c}(\mathbb{T}^{d}\times\mathbb{R}^{d})$. Here, $f_{0}$ is the part of the solution that is independent of $\varepsilon$, and $f_{1}$ corresponds to the first order effects in $\varepsilon$. Integrating this expansion in space, we have $\rho_{f_{\varepsilon}}=\rho_{0}+\varepsilon\rho_{1}+\mathcal{O}(\varepsilon^{2})$ where $\rho_{i}:=\int_{\mathbb{T}^{d}}f_{i}\,\mathrm{d}x$ for $i=\{0,1\}$. Substituting these expansions into (1.2) and letting $\varepsilon\rightarrow 0$ we immediately find $f_{0}=\mathcal{M}\rho_{0}$.
Assuming the boundedness from below of $\rho_{0}$, we now expand the power law as a binomial series to obtain
@@ 1.3 State of the art
In a series of papers [5, 7, 6], Bardos, Golse, Perthame and Sentis study the Rosseland approximation for the radiative transfer equation which is a nonlinear kinetic transport equation similar to (1.1). This equation models the transport of photons in a starlike medium. The term corresponding to the collision rate $\lambda(\rho):=\rho^{\alpha}$ in (1.1) is the opacity of the medium $\sigma(T)$ at the internal energy $T$. In [5, 7, 6], the authors derive the so-called Rosseland equation as a hydrodynamic limit of the radiative transfer equation. The Rosseland equation is a degenerate parabolic equation similar to (1.3) with $\alpha=1$ and a constant boundary value. Particularly [7] covers partially the case corresponding to $1<\alpha<0$ in (1.1) without requiring the monotonicity of $\sigma$. They employ Schauder’s fixed point argument, some compactness results and energy estimates to prove the well-posedness of the kinetic equation.
In [20], Dolbeault, Markovich, Oelz and Schmeiser study a different type of kinetic transport equation with a nonlinear relaxation term towards a generalised local Gibbs state. The equation is posed in $(x,v)\in(\mathbb{R}^{3}\times\mathbb{R}^{3})$ with a given confining potential. In the diffusive limit, they obtain drift-diffusion equations with a nonlinear diffusion term of porous medium type, corresponding to the case $\frac{2}{3}<\alpha<1$ in (1.3). The drift term arises from the confining potential. They prove the existence and uniquness of the solutions of the kinetic equation for initial data bounded by equilibrium distributions. The diffusion limit is obtained by using a div-curl lemma based compactness argument; thus, is not quantitative.
More recently, Anceschi and Zhu, in [2], study the Cauchy problem and diffusion asymptotics of a nonlinear kinetic Fokker-Planck type equation (1.5) where $\alpha\in[0,1]$, and $(x,v)\in\mathbb{T}^{d}\times\mathbb{R}^{d}$. They prove well-posedness and propagation of Maxwellian lower bounds of the solutions by using some results based on De Giorgi-Nash-Moser theory and the Harnack inequality. They derive quantitative diffusion asymptotics by combining entropic hypocoercivity, relative $\phi$-entropy, and barrier function methods. In this paper, we will follow their broad strategy for showing the quantitative diffusive limit. The key difference to our work is that the kinetic Fokker-Planck equation has smooth solutions.
@@
Hydrodynamic limits make up a large area of active research in kinetic theory, thanks to its connection to Hilbert’s sixth problem entitled "Mathematical Treatment of the Axioms of Physics" which concerns developing rigorously the limiting processes connecting the atomistic description of the matter with the laws of motion of continua. Therefore, the most classical problem in this area is justifying the limit from the Boltzmann equation to either Euler or Navier-Stokes equations. Breakthrough results are due to Bardos, Golse and Levermore in [3, 4] and then due to Golse and Saint-Raymond in [25]. For a good review in this direction we refer the reader to [34]. In [33], Saint-Raymond provided a complete derivation of the Navier–Stokes–Fourier equations from a BGK equation (1.4) when $\lambda$ is a constant. A major obstacle in resolving Hilbert’s sixth problem was the long-time justification of the Boltzmann equation. Recent works of Deng, Hani and Ma establishes a Boltzmann-Grad derivation valid up to the lifespan of smooth Boltzmann solutions in [17], and building on this and existing hydrodynamic-limit results, derive the compressible Euler and incompressible Navier–Stokes–Fourier systems on $\mathbb{T}^{d},\,d=\{2,3\}$ in [18]. These results represent the current state of the art on Hilbert’s sixth problem.
@@ 2.1 Main results
We present our main results with three theorems. In what follows, we denote the weighted $L^{2}$ norm of a function $f$ as $\|f\|_{L^{2}_{x,v}(\mathcal{M}^{-1})}:=\left(\int_{\mathbb{T}^{d}}\int_{\mathbb{R}^{d}}|f(t,x,v)|^{2}\mathcal{M}^{-1}(v)\,\mathrm{d}v\,\mathrm{d}x\right)^{\frac{1}{2}}$.
@@
Let $\varepsilon>0,A>1,\alpha\in\mathbb{R}$, and let $f_{\varepsilon}(t,x,v)$ be the solution to (1.2) with $\|f_{\varepsilon}\|_{L^{1}_{x,v}}=1$. Suppose that $\rho_{f_{\varepsilon}}(t,x)=\int_{\mathbb{R}^{d}}f_{\varepsilon}(t,x,v)\,\mathrm{d}v$ satisfies $A^{-1}\leq\rho_{f_{\varepsilon}}(t,x)\leq A$ for all $(t,x)\in[0,\infty)\times\mathbb{T}^{d}$. Then, there exists $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0}),t\geq 0$, we have
@@
$\gamma=\frac{1}{8A^{\alpha}(2+A^{2\alpha})}.$
@@
Remark 2.4.
The condition on the initial data simply requires that the initial data for the kinetic equation and the diffusion equation remain close in the limit $\varepsilon\rightarrow 0$. Taking $\varepsilon^{\prime}:=\varepsilon$, this is equivalent to a more classical well-preparedness condition on the initial data.
@@ 2.2 Preliminary results on hydrodynamic quantities
$\begin{split}\rho_{\varepsilon}(t,x)&:=\int f_{\varepsilon}(t,x,v)\,\mathrm{d}v,\\ j_{\varepsilon}(t,x)&:=\int f_{\varepsilon}(t,x,v)v\,\mathrm{d}v,\\ E_{\varepsilon}(t,x)&:=\int f_{\varepsilon}(t,x,v)(v\otimes v-I)\,\mathrm{d}v.\end{split}$ (2.2)
@@
Let $f_{\varepsilon}$ be the solution to (1.2) and $\rho_{\varepsilon},j_{\varepsilon},E_{\varepsilon}$ be the hydrodynamic quantities associated to $f_{\varepsilon}$ as defined in (2.2). Then,
@@
$\partial_{t}(\varepsilon\nabla_{x}\cdot(\rho_{\varepsilon}^{-\alpha}j_{\varepsilon}))$ $=-\alpha\nabla_{x}\cdot(\rho_{\varepsilon}^{-\alpha-1}j_{\varepsilon}(\nabla_{x}\cdot j_{\varepsilon}))-\nabla_{x}\cdot(\rho_{\varepsilon}^{-\alpha}\nabla_{x}\cdot E_{\varepsilon})-\nabla_{x}\cdot(\rho_{\varepsilon}^{-\alpha}\nabla_{x}\rho_{\varepsilon})-\frac{1}{\varepsilon}\nabla_{x}\cdot j_{\varepsilon}.$
@@
Similarly from equations (2.3), (2.4), we can see that we expect $\frac{1}{\varepsilon}\rho_{\varepsilon}^{\alpha}j_{\varepsilon}-\nabla_{x}\rho_{\varepsilon}\rightarrow 0,$ as $t\rightarrow\infty$.
@@ 3 Well-posedness
for $t\in\mathbb{R}_{+},x\in\mathbb{T}^{d},v\in\mathbb{R}^{d}$ and $\lambda:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$. We assume that $\lambda$ is a $C^{1}$ function satisfying the following bounds,
@@
$\partial_{t}r+v\cdot\nabla_{x}r+\lambda(\rho_{f})r$ $=\lambda(\rho_{f})(A-\rho_{f})\mathcal{M},$
@@
By the method of characteristics we may explicitly solve for $r$ and use $r(t=0)\geq 0$, $\rho_{f}\leq A$ to show that $r\geq 0$, and so $g\leq A\mathcal{M}$. Similarly, taking $r=g-A^{-1}\mathcal{M}$, we find that $A^{-1}\mathcal{M}\leq g$ and so $g\in\mathcal{V}$.
Let us now show that $\Gamma$ is a contraction on $\mathcal{V}$. To this end, we let $f_{1},f_{2}\in\mathcal{V}$, set $g_{1}=\Gamma(f_{1}),g_{2}=\Gamma(f_{2})$ and $w=g_{1}-g_{2}$. Then $w$ satisfies
$\partial_{t}w+v\cdot\nabla_{x}w$ $=\lambda(\rho_{f_{1}})\rho_{f_{1}}\mathcal{M}-\lambda(\rho_{f_{1}})g_{1}-\lambda(\rho_{f_{2}})\rho_{f_{2}}\mathcal{M}+\lambda(\rho_{f_{2}})g_{2}$
$=\left(\lambda(\rho_{f_{1}})\rho_{f_{1}}-\lambda(\rho_{f_{2}})\rho_{f_{2}}\right)\mathcal{M}-g_{2}\bigl(\lambda(\rho_{f_{1}})-\lambda(\rho_{f_{2}})\bigr)-\lambda(\rho_{f_{1}})w.$
Collecting all terms involving $w$ on the left-hand-side, we find
@@ 4 Long-time behaviour
In this section, we show that the solution to Equation (1.2) converges to equilibrium as $t\rightarrow\infty$, in the space $L_{x,v}^{2}(\mathcal{M}^{-1}):=\left\{g\,:\,\int_{\mathbb{T}^{d}}\int_{\mathbb{R}^{d}}\frac{|g(t,x,v)|^{2}}{\mathcal{M}(v)}\,\mathrm{d}v\,\mathrm{d}x<\infty\right\}$. We achieve this by adapting the hypocoercivity techniques developed in [21]. Particularly, we have to pay attention to ensure that the nonlinear terms do not cause us problems.
@@ 4.1 $L^{2}$-hypocoercivity à la Dolbeault-Mouhot-Schmeiser
We consider $u(x),\,x\in\mathbb{T}^{d}$ and define its Fourier transform $\hat{u}(k):=\int_{\mathbb{T}^{d}}u(x)e^{-2\pi ikx}\,\mathrm{d}x$. We also define the operator $(I-\Delta_{x})^{-1}$ for $x\in\mathbb{T}^{d}$ by $\mathscr{F}((I-\Delta_{x})^{-1}u(x)):=\frac{1}{1+4\pi^{2}|k|^{2}}\hat{u}(k)$.
@@
$\|(I-\Delta_{x})^{-1}\partial_{x_{i}}u\|_{L^{2}_{x}}^{2}$ $\leq\|u-\bar{u}\|_{L^{2}_{x}}^{2}\leq\|u\|_{L^{2}_{x}}^{2},$ (4.2)
$\|(I-\Delta_{x})^{-1/2}\nabla_{x}u\|_{L^{2}_{x}}^{2}$ $\geq C\|u-\bar{u}\|_{L^{2}_{x}}^{2},$ (4.3)
where $\bar{u}:=\int_{\mathbb{T}^{d}}u(x)\,\mathrm{d}x$ and $C=4\pi^{2}/(1+4\pi^{2})$.
@@
Notice that the operator $(I-\Delta_{x})^{-1}$ commutes with taking derivatives of $u$ in $x$, i.e., $\nabla_{x}((I-\Delta_{x})^{-1}u)=(I-\Delta_{x})^{-1}\nabla_{x}u$. It is also positive and self adjoint and so has a square root which is given by $\mathscr{F}((I-\Delta_{x})^{-1/2}u)=\frac{1}{(1+4\pi^{2}|k|^{2})^{1/2}}\hat{u}$. By Plancherel theorem, we have $\|u\|^{2}_{L^{2}_{x}}=\sum_{k\in\mathbb{Z}^{d}}|\hat{u}(k)|$. Moreover, for the zero-th Fourier mode, i.e., $k=0$, we have that $\hat{u}(0)=\int_{\mathbb{T}^{d}}u(x)\,\mathrm{d}x=\bar{u}$. This means
$\|u-\bar{u}\|^{2}_{L^{2}_{x}}=\sum_{k\in\mathbb{Z}^{d},k\neq 0}|\hat{u}(k)|^{2}\leq\sum_{k\in\mathbb{Z}^{d}}|\hat{u}(k)|^{2}=\|u\|^{2}_{L^{2}_{x}}.$
The rest follows by direct computation. Then, we have
$\|(I-\Delta_{x})^{-1}\partial^{2}_{x_{i},x_{j}}u\|_{L^{2}_{x}}^{2}$ $=\sum_{k\in\mathbb{Z}^{d}}\left|\frac{-4\pi^{2}k_{i}k_{j}}{1+4\pi^{2}|k|^{2}}\hat{u}(k)\right|^{2}\leq\sum_{k\in\mathbb{Z}^{d}}|\hat{u}(k)|^{2}=\|u\|_{L^{2}_{x}}^{2},$
$\|(I-\Delta_{x})^{-1}\partial_{x_{i}}u\|_{L^{2}_{x}}^{2}$ $=\sum_{k\in\mathbb{Z}^{d}}\left|\frac{2\pi ik_{i}}{1+4\pi^{2}|k|^{2}}\hat{u}(k)\right|^{2}\leq\sum_{k\in\mathbb{Z}^{d},k\neq 0}|\hat{u}(k)|^{2}=\|u-\bar{u}\|_{L^{2}_{x}}^{2}\leq\|u\|_{L^{2}_{x}}^{2},$
$\|(I-\Delta_{x})^{-1/2}\nabla_{x}u\|_{L^{2}_{x}}^{2}$ $=\sum_{k\in\mathbb{Z}^{d}}\frac{4\pi^{2}|k|^{2}}{1+4\pi^{2}|k|^{2}}|\hat{u}(k)|^{2}\geq C\sum_{k\in\mathbb{Z}^{d},k\neq 0}|\hat{u}(k)|^{2}=C\|u-\bar{u}\|_{L^{2}_{x}}^{2},$
where $C=4\pi^{2}/(1+4\pi^{2})$. ∎
@@
First we look at the dissipation of the $L^{2}$ norm. Occasionally we use the shorter notations $f_{\varepsilon}(v)=f_{\varepsilon}(t,x,v)$ or $f_{\varepsilon}(t)=f_{\varepsilon}(t,x,v)$. We first note that,
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\int_{\mathbb{T}^{d}}\int_{\mathbb{R}^{d}}\frac{(f_{\varepsilon}-\mathcal{M})^{2}}{\mathcal{M}}\,\mathrm{d}v\,\mathrm{d}x=2\int_{\mathbb{T}^{d}}\int_{\mathbb{R}^{d}}\frac{f_{\varepsilon}-\mathcal{M}}{\mathcal{M}}\partial_{t}f_{\varepsilon}\,\mathrm{d}v\,\mathrm{d}x=2\int_{\mathbb{T}^{d}}\int_{\mathbb{R}^{d}}\frac{f_{\varepsilon}}{\mathcal{M}}\partial_{t}f_{\varepsilon}\,\mathrm{d}v\,\mathrm{d}x,$
where the second term vanishes due to the conservation of mass. We then have,
@@
$=-\frac{1}{\varepsilon^{2}}\int_{\mathbb{T}^{d}}\rho_{\varepsilon}^{\alpha}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left(\left(\frac{f_{\varepsilon}(u)}{\mathcal{M}(u)}-\frac{f_{\varepsilon}(v)}{\mathcal{M}(v)}\right)^{2}+\frac{f_{\varepsilon}^{2}(u)}{\mathcal{M}^{2}(u)}-\frac{f_{\varepsilon}^{2}(v)}{\mathcal{M}^{2}(v)}\right)\mathcal{M}(u)\mathcal{M}(v)\,\mathrm{d}u\,\mathrm{d}v\,\mathrm{d}x$